# CDKFormer: Contextual Deviation Knowledge-Based Transformer for Long-Tail Trajectory Prediction

### v1

@@ V-D1 Ablation Study on Input Deviation Feature
TABLE IV: Ablation study on input deviation feature. The results are presented in minADE6/minFDE6.
@@
As shown in Table IV, the results of the ablation study indicate that incorporating both individual deviation and group deviation features significantly enhances the model performance, particularly in challenging scenarios. A baseline model with no deviation feature reaches a minADE6 of 0.78 m, 1.38 m and 1.50 m on all, top 10% and top 5% samples, respectively. When only individual deviation is included, the model demonstrates certain improvements. On the top 10% tailed samples, where the minADE6 is reduced to 1.35 m from 1.38 m, and the minFDE6 is 2.62 m. Similarly, including only group deviation also leads to moderate performance gains. The best overall performance is achieved when both individual deviation and group deviation are used together. The model achieves a minADE6 of 0.75 m and a minFDE6 of 1.42 m. On tailed samples, we also observe a notable improvement over the baseline model, with a 3.62% and 4.00% reduction of minADE6 across the top 10% and 5% samples. This highlights the importance of integrating both individual and group deviation features to improve long-tail trajectory prediction.
V-D2 Ablation Study on Query Design
In this ablation study, we evaluate the effect of different decoding queries on the performance of the CDKFormer decoder, as shown in Table V. The loss of one query is abandoned if the corresponding query is removed. The deviation feature is always retained in the decoder.
TABLE V: Ablation study on query design. The results are presented in minADE6/minFDE6.
@@ V-D2 Ablation Study on Query Design
TABLE VI: Ablation study on multistream cross-attention block structure
All Top 10% Top 5%
Stream sequence #Layers minADE6 (m) minFDE6 (m) minADE6 (m) minFDE6 (m) minADE6 (m) minFDE6 (m)
@@
The results indicate that using only the mode query or the regular future query leads to a performance drop, particularly on the tailed samples. A marked improvement is observed when the mode query and regular future query are combined, yielding a minFDE6 of 1.42 m and 2.60 m on all samples and top 10% samples, respectively. Pairing the mode query with the tail future query achieves a minADE6 of 0.77 m and 1.48 m on all samples and top 5% samples. However, this configuration performs slightly worse than the mode and regular future query combination. The optimal performance is attained when all three queries are utilized together. A 0.77% and 0.68% reduction in minFDE6 on top 10% and top 5% samples are observed, compared to the model with only the mode and regular future query. These findings highlight the importance of incorporating dual queries to effectively capture temporal patterns and enhance robustness in long-tail trajectory prediction scenarios.
V-D3 Ablation Study on Multistream Cross-Attention Block Structure
@@
Figure 8: Visualization of CDKFormer multimodal trajectory prediction results. Vehicles are represented by bounding boxes in dark blue. Historical and ground-truth trajectories are illustrated with light blue and orange solid lines, respectively. The predicted multimodal future trajectories are shown by yellow dash lines.
V-E Qualitative Analysis
@@
To evaluate the reliability of the proposed model, particularly in challenging conditions, we present visualizations of the trajectory prediction results in three representative long-tail scenarios. Fig. 8(a) depicts the target vehicle executing a left turn at an intersection. The model’s multimodal predictions accurately align with the lane segments and capture the left turn maneuver, accounting for various possible speeds. Fig. 8(b) presents a more complex scenario in which the target vehicle proceeds straight through the intersection. The proposed CDKFormer successfully predicts the trajectory required to go straight and also generates alternative trajectories for a possible right turn, demonstrating its ability to handle multiple potential behaviors. Fig. 8(c) illustrates a scenario involving a long-range trajectory where the target vehicle executes a left turn, with the intention of the target vehicle already shown from its historical trajectory. CDKFormer accurately predicts the future path, showcasing its effectiveness in leveraging past motion and deviation information for accurate forecasting.
VI Conclusions
@@
In this study, we introduce CDKFormer, a novel framework tailored for long-tail trajectory prediction. Our approach addresses the challenges posed by long-tailed distributions in trajectory prediction tasks, ensuring robust performance across diverse scenarios. This study begins with a comprehensive analysis of long-tail characteristics of a large-scale trajectory prediction dataset, from which we derive features that effectively characterize long-tailed samples. Leveraging the extracted features, we propose the Contextual Deviation Knowledge-based Transformer (CDKFormer) model. We design a scene context encoding module and a deviation feature fusion module composed of Transformer encoder layers to integrate scene contextual information and obtain a comprehensive representation of the driving environment. Subsequently, a dual query-based decoder is developed. Employing a multistream decoder block, we leverage a mode query and dual future queries to decode heterogeneous scene deviation features. The dual queries, including regular and tail future queries, are specifically designed to encapsulate both normal-state and tail-state information. These queries are then integrated into the standard scene query, enabling subsequent refinement and multimodal trajectory generation.
@@
We evaluate the proposed model using the Argoverse 2 Motion Forecasting dataset, where CDKFormer achieves SOTA performance across multiple evaluation metrics, confirming its effectiveness in predicting future trajectories under long-tailed conditions. Ablation studies further substantiate the contributions of each component, highlighting their individual and collective impact on overall performance. Collectively, our method provides a robust framework for understanding long-tailed scenarios and introduces a new perspective on trajectory prediction models in rare and challenging scenarios.
One possible direction for future research is to expand deviation modeling to account for multiple interacting agents within complex traffic environments. While the present study primarily centers on deviation modeling for the target vehicle, exploring the relationship between the deviation states of surrounding agents and subsequently using it for multiagent prediction could significantly improve the model performance in real-world settings. Another promising direction is to account for heterogeneity among surrounding agents. While the current CDKFormer can handle a variety of traffic agents as the inputs, explicitly considering the types of surrounding agents and modeling their heterogeneity in long-tail trajectory prediction could significantly enhance the model’s performance. This approach would better capture the diverse behaviors of pedestrians, cyclists, vehicles, etc., potentially benefiting the robustness of autonomous systems in complex and real-world scenarios.

### v2

@@ 5.4.1 Ablation Study on Input Deviation Feature
Table 7: Ablation study on input deviation feature. Results are shown on Argoverse 2 motion forecast dataset and presented in minADE6/minFDE6.
@@
As shown in Table 7, the results of the ablation study indicate that incorporating both individual deviation and group deviation features significantly enhances the model performance, particularly in challenging scenarios. A baseline model with no deviation feature reaches a minADE6 of 0.78 m, 1.38 m and 1.50 m on all, top 10% and top 5% samples, respectively. When only individual deviation is included, the model demonstrates certain improvements. On the top 10% tail samples, where the minADE6 is reduced to 1.35 m from 1.38 m, and the minFDE6 is 2.62 m. Similarly, including only group deviation also leads to moderate performance gains. The best overall performance is achieved when both individual deviation and group deviation are used together. The model achieves a minADE6 of 0.75 m and a minFDE6 of 1.42 m. On tail samples, we also observe a notable improvement over the baseline model, with a 3.62% and 4.00% reduction of minADE6 across the top 10% and 5% samples. This highlights the importance of integrating both individual and group deviation features to improve long-tail trajectory prediction.
5.4.2 Ablation Study on Query Design
In this ablation study, we evaluate the effect of different decoding queries on the performance of the CDKFormer decoder, as shown in Table 8. The loss of one query is abandoned if the corresponding query is removed. The deviation feature is always retained in the decoder.
Table 8: Ablation study on query design. Results are shown on Argoverse 2 motion forecast dataset and presented in minADE6/minFDE6.
@@ 5.4.2 Ablation Study on Query Design
The results indicate that using only the mode query or the regular future query leads to a performance drop, particularly on the tail samples. A marked improvement is observed when the mode query and regular future query are combined, yielding a minFDE6 of 1.42 m and 2.60 m on all samples and top 10% samples, respectively. Pairing the mode query with the tail future query achieves a minADE6 of 0.77 m and 1.48 m on all samples and top 5% samples. However, this configuration performs slightly worse than the mode and regular future query combination. The optimal performance is attained when all three queries are utilized together. A 0.77% and 0.68% reduction in minFDE6 on top 10% and top 5% samples are observed, compared to the model with only the mode and regular future query. These findings highlight the importance of incorporating dual queries to effectively capture temporal patterns and enhance robustness in long-tail trajectory prediction scenarios.
5.4.3 Ablation Study on MultiStream Cross-Attention Block Structure
Table 9: Ablation study on multistream cross-attention block structure. Results are shown on Argoverse 2 motion forecast dataset and presented in minADE6/minFDE6.
Stream sequence #Layers All Top 10% Top 5%
@@ 5.4.3 Ablation Study on MultiStream Cross-Attention Block Structure
Figure 13: Visual comparison of multimodal trajectory prediction results. Vehicles are represented by bounding boxes in dark blue. Historical and ground-truth trajectories are illustrated with light blue and orange solid lines, respectively. The top row displays the results from our proposed CDKFormer, while the bottom row shows the corresponding predictions from QCNet. The corresponding difficulty score $S_{d}$ and rarity score $S_{r}$ are provided for each scenario.
5.5 Qualitative Analysis
To provide a tangible assessment of the model’s performance in challenging conditions, Figure 13 presents a visual comparison between CDKFormer and QCNet across five representative long-tail scenarios. This highlights that CDKFormer consistently generates more accurate and plausible multimodal predictions.
@@ 5.5 Qualitative Analysis
Figure 13(a) presents a complex scenario in which the target vehicle proceeds straight through the intersection and turn left at the end of the prediction horizon. The proposed CDKFormer successfully predicts the trajectory required to turn left and also generates alternative trajectories for a possible straight move, demonstrating its ability to handle multiple potential behaviors. A similar scenario is also shown in Figure 13(c), which depicts the target vehicle executing a left turn at an intersection. CDKFormer’s multimodal predictions accurately align with the lane segments and capture the left-turn maneuver, accounting for various possible speeds, while QCNet produces highly concentrated and implausible trajectories. In challenging and rule-violating scenarios, our model also demonstrates reliability. For example, when a vehicle enters from outside the drivable area to make an unexpected right turn (Figure 13(b)), the predicted trajectories of CDKFormer align more closely with the ground truth than those of QCNet. In the challenging pull-over scenario in Figure 13(d), both models struggle to precisely match the ground-truth. However, CDKFormer’s best predicted mode terminates closer to the vehicle’s final stopping position. This demonstrates a more accurate understanding of the vehicle’s final goal, critical for safety in downstream planning. A similar advantage is observed in Figure 13(e), where CDKFormer’s predicted trajectories are more tightly clustered around the ground-truth path and are more constrained and realistic.
6 Conclusions
@@
In this study, we introduce CDKFormer, a novel framework tailored for long-tail trajectory prediction. Our approach addresses the challenges posed by long-tailed distributions in trajectory prediction tasks, ensuring robust performance across diverse scenarios. This study begins with a comprehensive analysis of the long-tail characteristics of a large-scale trajectory prediction dataset, from which we derive features that effectively characterize long-tail samples. Leveraging extracted features, we propose a contextual deviation knowledge-based Transformer (CDKFormer) model. We design a scene context encoding module and a deviation feature fusion module composed of Transformer encoder layers to integrate scene contextual information and obtain a comprehensive representation of the driving environment. Subsequently, a dual query-based decoder is developed. Employing a multistream decoder block, we leverage a mode query and dual future queries to decode heterogeneous scene deviation features. The dual queries, including regular and tail future queries, are specifically designed to encapsulate both normal-state and tail-state information. These queries are then integrated into the standard scene query, enabling subsequent refinement and multimodal trajectory generation.
@@
We evaluate the proposed model using the Argoverse 2 motion forecasting dataset and inD dataset, where CDKFormer achieves SOTA performance across multiple evaluation metrics, confirming its effectiveness in predicting future trajectories under long-tailed conditions. Ablation studies further substantiate the contributions of each component, highlighting their individual and collective impact on overall performance. Collectively, our method provides a robust framework for understanding long-tailed scenarios and introduces a new perspective on trajectory prediction models in rare and challenging scenarios.
One possible direction for future research is to incorporate map-based deviation modeling. This would involve explicitly identifying long-tail scenarios related to an agent’s non-compliance with road semantics, such as trajectories that cross solid lane boundaries, cut across intersections improperly, or otherwise deviate from the expected road topology. Integrating these map-aware deviation signals could provide the model with a contextually grounded understanding of what constitutes a rare event. Furthermore, understanding the causes of long-tail prediction failures remains a significant challenge for the field. This study offers initial insights by correlating long-tail scenarios with motion-based deviations. Nevertheless, a deeper causal analysis, moving from what fails to why it fails, is essential for building robust and reliable prediction models.
Acknowledgement(s)
The research is supported by National Natural Science Foundation of China 52325209, 52272420, Tsinghua University-Mercedes Benz Joint Institute for Sustainable Mobility, and Tsinghua-Toyota Joint Research Institute Inter-disciplinary Program.